This coming summer, we are meeting up with some family members at Family Fiesta Amusement Park. The ticket price for children is $17 and for adults the ticket price is $31. The total price for all 20 of us to enter the park is $438.
I tried A+C=20, A31+C17=438, With that I tried wlimination method and got an anwser of 0.83. how do I solve this one Please????

Answer by stanbon(75887)

This coming summer, we are meeting up with some family members at Family Fiesta Amusement Park. The ticket price for children is $17 and for adults the ticket price is $31. The total price for all 20 of us to enter the park is $438.
----
People Eq.::: a + c = 20
Value Eq.:: 31a+17c = 438
----
Multiply thru People Eq. by 31 to get:
31a + 31c = 31*20
----
Subtract from the Value Eq and solve for "c":
-14c = 438-31*20
-14c = -182
---
c = 13 (# of child tickets sold)
---
Since a + c = 20, a = 7 (# of adult tickets sold)
===================
Cheers,
Stan H.